Adovbush:

A box contains 8 rings, 10 necklaces, and 4 pins. If Anya selects 1 piece at random, does NOT replace it and then picks another, what is the probability that she selected 2 rings?

SmokeyBrown:

This one's actually pretty similar to the last one we did. So, to start, picking the first ring would have a 8/22 chance... Can you figure out what the chance of picking the second ring would be (assuming no replacement)? And then, you'll recall, the total probability of the outcome is the probability of both events multiplied together

Adovbush:

7/21?

snowflake0531:

First add the total amount of... things in the box 8+10+4 = 22 So the first time she picks out a ring, she has 8/22 chance of getting it because there are 8 rings in all, but then it says 'does not replace and chooses another piece, which means there are 21 pieces left and 7 rings so multiply \[ \frac {8} {22} * \frac {7} {21} \]

Adovbush:

4/33 is the answer?

snowflake0531:

ye

Adovbush:

ight thanks both yall